OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (gabbyalicorn):

So we have the equation 5(8+7)=5x+(5)(7) Step 1: Simplify both sides of the equation. 75=5x+35 Step 2: Flip the equation. 5x+35=75 Step 3: Subtract 35 from both sides. 5x+35−35=75−35 5x=40 Step 4: Divide both sides by 5. 5x5=405 x=8 Answer: x=8

OpenStudy (anonymous):

The answer would be 8. Step 1: Simplify both sides of the equation. 75=5x+35 Step 2: Flip the equation. 5x+35=75 Step 3: Subtract 35 from both sides. 5x+35−35=75−35 5x=40 Step 4: Divide both sides by 5. 5x/5=40/5 x=8 (mathpapa.com is a really good site for solving these equations)

OpenStudy (anonymous):

Thank you!!!